OpenStudy (anonymous):

A trapezoid has an area given by A = (b 1 + b 2)h, where b 1 is the length of one base, b 2 is the length of the other base, and h is the height. If the trapezoid has an area of 14 square units, solve for b 1 in terms of the other variables.

jigglypuff314 (jigglypuff314):

A = (b1 + b2)h first distribute the h A = b1 h + b2 h then subtract b2 h from both sides A - b2 h = b1 h then divide both sides by h b1 = A/h - b2 plug in 14 for A b1 = ....

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

what does b1 =

jigglypuff314 (jigglypuff314):

b1 = A/h - b2 then to get the answer that the question wants you to get, plug in 14 for A

OpenStudy (anonymous):

Ok. Thanks